# Tapping into the Black Box: Uncovering Aligned Representations in Pretrained Neural Networks

### v1

@@ 3.3 Path space and the linear nature of ReLU networks
Consider $\mathcal{P}_{0}$ i.e. the set of paths through layers $0,\ldots,L$ and $p=(p_{0},p_{1},\ldots,p_{L})$. Let $p|_{k}:=(p_{k},\ldots,p_{L})$, i.e. the (unique) sub-path of $p$ starting at layer $k$. We will abuse the notation and write $\tau_{p}$ instead of $\tau_{p|_{1}}$. With that remark we can expand $f_{\tau}(x)$ as:
@@
Again, we call $\mathbb{R}^{P_{0}}$ a tensor product as $\mathbb{R}^{P_{0}}\cong\bigotimes_{\ell=0}^{L}\mathbb{R}^{d_{\ell}}\cong(\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}})\otimes\mathbb{R}^{d_{0}}$. Additionally, we refer to $\mathbb{R}^{P_{0}}$ as the path space, as this is the space of all the neural paths in the network.
@@
$\Omega((y,M)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (17)
@@
Notice that the $\tau$-pullback defined earlier (Equation 15) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 16 and 19, we have:
@@
Note 3.3 The equation 20 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$, and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$. This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$. Moreover, by Equation 20, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$.
Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}$, highlighting the importance of the path-representation.
4 Excitation pullback and the implicit linear model
Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance. In particular, weight pruning techniques demonstrate that many weights in trained networks can be removed without significantly degrading accuracy, implying the presence of sparse, efficient subnetworks within dense architectures (Han et al., 2015). Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation. Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure even before convergence. Further motivation is provided in Section 6.1.
Let $X\subset\mathbb{R}^{d_{0}}$ be a fixed, compact dataset (e.g. $[0,1]^{d_{0}}$ in case of images) and let $g,h:\mathbb{R}^{d_{0}}\to\mathbb{R}^{n}$ be two functions. Define:
$\rho(x):=\operatorname{corr}(g(x),h(x))$
where $\operatorname{corr}$ denotes the Pearson correlation coefficient between the vectors $g(x)$ and $h(x)$. We say that $g$ and $h$ are positively correlated (in $X$) if $\rho(x)$ is positive for almost every $x\in X$, with respect to the Lebesgue measure. Similarly, we’ll call the correlation strong if $\rho(x)$ is high (e.g., greater than 0.7, although the exact threshold is not important) almost everywhere in $X$. Our core hypothesis is as follows:
Hypothesis 1 (Early emergence of stable excitation patterns).
There exists a binary tensor field $\gamma:\mathbb{R}^{d_{0}}\to\{0,1\}^{\mathcal{P}1}$ and an early training time $t\gamma\ll T$ such that for every $t>t_{\gamma}$, $\gamma$ is positively correlated with $\tilde{G}^{(t)}$ (the gating tensor field at time $t$).
Moreover, we assume that the network encodes the field $\gamma$ in its pre-activations. That is, there exists a gating function $\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}}\subset\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ such that:
  • $\Gamma$ is strongly positively correlated with the pre-activation function $Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$,
  • the induced rank-1 tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to[0,1]^{\mathcal{P}1}$ is strongly positively correlated with $\gamma$.
Comments and intuition:
  1. 1.
    We assume that the field $\gamma$ has binary weights and therefore acts as a path filter, just like $\tilde{G}$; however, it is not necessarily a rank-1 tensor, unlike $\tilde{G}$.
  2. 2.
    The field $\tilde{\Gamma}$ is composed of rank-1 tensors and thus can be seen as a uniform rank-1 approximation of $\gamma$.
  3. 3.
    Intuitively, $\gamma$ contains paths that pass through highly excited neurons (those with strong pre-activations).
  4. 4.
    The field $\tilde{\Gamma}$ provides a smoother approximation to $\gamma$ than any hard-gated field $\tilde{G}^{(t)}$, since it performs a form of soft thresholding; in particular, it suppresses barely active paths (i.e. those passing through weakly activated neurons) and partially recovers paths that are nearly active (i.e. those interrupted by a single weak neuron, but otherwise strong). This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and more aligned with the coarser structure of the decision boundary.
  5. 5.
@@ 4 Excitation pullback and the implicit linear model
    The function $\Gamma$ can be approximated as $\sigma\circ Z$, where $\sigma$ is a sigmoid-like nonlinearity applied element-wise, e.g. sigmoid, softsign (with properly adjusted range), normal cumulative distribution function (normal CDF), etc. In general, $\sigma$ should be neuron-specific or at least layer-specific, since it depends on the pre-activation distribution, which may be different for each neuron. In practice, thanks to the BatchNorm layers (Ioffe and Szegedy, 2015), a shared $\sigma$ across all neurons often yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters).
  6. 6.
    $\tilde{\Gamma}$ approximates the direction of $\gamma$ (in the sense of cosine similarity) in $\mathbb{R}^{P_{1}}$, rather than its exact $L_{2}$-magnitude, since $\tilde{\Gamma}_{p}\ll 1$ for most paths $p$.
We refer to $\Gamma$ as the excitation function of the network and to $v_{\Gamma}$ as the excitation pullback.
Note 4.1 If the Hypothesis 1 holds, then, starting from time $t_{\gamma}$, the training dynamics of the network closely resembles the dynamics of a linear model $f_{\gamma}$ with a fixed feature map $\phi_{\gamma}$ (see Equation 18), i.e. a kernel machine operating under the feature noise introduced by $\phi_{\tilde{G}^{(t)}}$. In particular, the decision boundary of that implicit linear model $f_{\gamma}$ can be closely approximated by the decision boundary of $f_{\Gamma}$, the latter being directly visualizable by the excitation pullback $v_{\Gamma}$ (see Section 5).
Thus, the postulated $\gamma(x)$, potentially being a higher-rank tensor, would grasp the essence of what stays the same across different rank-1 neural responses $G^{(t)}(x)$ to the same input stimuli $x$ at time $t$ during network training.

### v3

@@ 3.3 Path space and the linear nature of ReLU networks
Consider $\mathcal{P}_{0}$, the set of all paths through layers $0,\ldots,L$ and $p=(p_{0},p_{1},\ldots,p_{L})\in\mathcal{P}_{0}$. Let $p|_{k}:=(p_{k},\ldots,p_{L})$ denote the (unique) sub-path of $p$ starting at layer $k$. We will abuse the notation and write $\tau_{p}$ instead of $\tau_{p|_{1}}$. With that remark we can expand $f_{\tau}(x)$ as:
@@
Again, we call $\mathbb{R}^{P_{0}}$ a tensor product as $\mathbb{R}^{P_{0}}\cong\bigotimes_{\ell=0}^{L}\mathbb{R}^{d_{\ell}}\cong(\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}})\otimes\mathbb{R}^{d_{0}}$. Additionally, we refer to $\mathbb{R}^{P_{0}}$ as the path space, as this is the space of all neural paths in the network.
@@
$\Omega((y,\theta)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
@@
Notice that the $\tau$-pullback defined earlier (Equation 14) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 15 and 18, we have:
@@
Note 3.3 The equation 19 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$ and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$. This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$. Moreover, by Equation 19, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$. Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}\in\mathbb{R}^{d_{L}}$, hinting at the importance of this representation.
4 Excitation pullbacks and the inherent gradient noise
4.1 On the inherent gradient noise
Despite the following representation property (see Equations 7 and 8):
$f(x)$ $=\langle v_{G}(x),x\rangle=v_{G}^{\top}(x)x$
$\nabla f(x)$ $=v_{G}^{\top}(x)|_{\mathbb{R}^{d_{0}^{*}}}$
the gradients of ReLU networks are notoriously hard to interpret (see Figure 2). This may suggest that those models rely on some strange predictive patterns that are meaningless to humans. But it may also mean that gradients are contaminated by an inherent noise that obfuscates a more regular underlying decision boundary. In this section, we advocate the latter view.
Indeed, by Equation 6 the pullback vector $v_{G}(x)$, and hence also the gradient $\nabla f(x)$, depend solely on the network parameters $(y,\theta)$ and the gating pattern $G(x)$. This means that $v_{G}(x)$ looses a lot of information about $x$, in particular about the pre-activation values $z_{\ell}(x)$. This loss of information might be important for the training process, as the network is less likely to overfit, but it is problematic for interpretability during inference.
To highlight the potential problems, suppose that $x\in\mathbb{R}^{d_{0}}$ lies on the boundary of two linear regions of $M$. Let $x_{a}\sim x$ and $x_{b}\sim x$ be two nearby points lying on the opposite sides of the boundary, i.e. each in a different linear region. By continuity of $f$, we have
$f(x_{a})\approx f(x)\approx f(x_{b})$
which, by Equation 7, translates to
$\langle v_{G}(x_{a}),x_{a}\rangle\approx\langle v_{G}(x),x\rangle\approx\langle v_{G}(x_{b}),x_{b}\rangle.$
By properties of the dot product we get
$\langle v_{G}(x_{a}),x\rangle\approx\langle v_{G}(x_{b}),x\rangle\implies\langle v_{G}(x_{a})-v_{G}(x_{b}),x\rangle\approx 0$ (20)
This means that, in general, the pullback vectors $v_{G}(x_{a})$ and $v_{G}(x_{b})$ may differ significantly as long as their difference remains orthogonal to $x$, i.e. it can be any point from the the entire hyperplane orthogonal to $x$. In standard training there is little incentive to align these orthogonal directions as they hardly affect the network’s value $f(x)$ (incidentally, such incentive may be provided by adversarial training, see Section 7.1).
Note 4.1 Intuitively, $v_{G}(x)$ may contain locally irrelevant directions from neurons that are spuriously active, i.e. barely above the activation threshold but not corresponding to any salient feature in $x$. On the other hand, some important neurons may be inactive, either due to the local noise or because $x$ is a negative example and does not contain many features relevant to the chosen class $c$ for which the gradient is computed.
4.2 Introducing excitation pullbacks
The considerations from previous section suggest that we should look beyond the hard activation gates $G(x)$ to properly pull the action of $y^{\top}$ back to $\mathbb{R}^{d_{0}}$ - in a way that is robust to hyper-local noise. A natural idea is to perform soft gating, i.e. compose the pre-activations $z_{\ell}$ with some sigmoid-like step function $\sigma_{\ell}:\mathbb{R}^{d_{\ell}}\to[0,1]^{d_{\ell}}$:
$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}}$ (21)
@@ 4.2 Introducing excitation pullbacks
In general, $\sigma_{\ell}$ is neuron-specific, since the pre-activation distribution may be different for each neuron. In practice, thanks to the BatchNorm layers, a shared $\sigma$ across all neurons yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters). In the paper we set the global element-wise $\sigma_{\ell}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}:\mathbb{R}\to[0,1]$, but other options (e.g. rescaled softsign or Normal CFD) lead to very similar results - as long as the general sigmoid shape of the global $\sigma$ is preserved, see Section 6.2.
Because $\sigma_{\ell}$ can be understood as measuring the excitation of a neuron we may accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field. Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to compute the right pre-activations.
Note 4.2 The tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ (see Equation 12) suppresses barely active paths, i.e. those passing through weakly activated neurons; and partially recovers paths that are nearly active (i.e. those interrupted by a single weak neuron, but otherwise strong). This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and probably more aligned with the coarser structure of the decision boundary. In a sense, $v_{\Gamma}$ is dominated by atoms corresponding to highly excited paths, as it quickly suppresses the less-excited paths (see Equation 10).
5 On the path stability and its potential significance
5.1 The path stability hypothesis
Because of the relative stability of $\phi_{\tilde{G}^{(t)}}$ (Note 3.3), it’s tempting to assume that the gating function $G^{(t)}$ gets fixed early on during training. We consider this assumption too optimistic as the gates probably do change simply because of the gradient noise we’ve hinted at before. Notice, however, that $\tilde{G}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ maps inputs to binary rank-1 tensors, meaning that it has limited path-filtering capability (Note 3.1). But perhaps there are more refined atom filters (corresponding to higher-rank binary tensors) that do get fixed early on during training? We motivate this assumption below.
Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance. In particular, weight pruning techniques demonstrate that many weights in trained networks can be removed without significantly degrading accuracy, implying the presence of sparse, efficient subnetworks within dense architectures (Han et al., 2015). Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation. Furthermore, in (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF), which are essentially our gating functions (Equation 3), and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”. Specifically, they introduce Deep Gated Networks (DGN) and provide empirical evidence that gate adaptation is the key to generalization, most notably, that “the winning lottery is in the gating pattern”  (Lakshminarayanan and Singh, 2020). Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure long before convergence.
Recall that $f^{(t)}(x)=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$ and $f^{(t)}_{\tau}(x)=\langle\omega^{(t)},\phi_{\tau}(x)\rangle$. Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f(X)=(f(x_{i}))_{i}$ be the vector of all the values of $f$. We say that two vectors $a,b\in\mathbb{R}^{n}$ are strongly positively correlated if the Pearson correlation coefficient $\rho(a,b)$ is high (e.g., greater than 0.9, although the exact threshold value is not important). We denote this relation as $a\approx_{+}b$. We now state our core hypothesis:
Hypothesis 1 (Early emergence of stable paths).
There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that for every $t>t_{\pi}$ we have $f^{(t)}(X)\approx_{+}f^{(t)}_{\pi}(X)$.
Additionally, we postulate that the network encodes the tensor field $\pi$ in its pre-activation function $Z$. That is, there exists a choice of $(\sigma_{\ell})_{\ell}$ (as in Equation 21) such that for every $x\in X$ we have $\tilde{\Gamma}(x)\approx_{+}\pi(x)$.
We refer to the above assumption as the path stability condition and to any such $\pi$ as path tensor feature.
Intuition:
The hypothesis states that, even though the gating tensors $\tilde{G}^{(t)}(x)$ change during training, there exists a time-independent gating tensor $\pi(x)$ that, starting from time $t_{\pi}$, largely determines network’s decision $f^{(t)}(x)$ at $x$. In this view, the variability of tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 tensors; the network, trying to approximate $\pi(x)$, learns it implicitly in it’s pre-activation pattern $Z(x)$ in the form of the set of highly-excited paths. The latter can be feasibly directionally approximated by the rank-1 non-binary tensor $\tilde{\Gamma}(x)$.
5.2 The implicit kernel machine and the faithfulness of excitation pullbacks
Recall that $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ (Equation 16) encodes the network’s parameters at time $t$ and that:
$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$
$f^{(t)}_{\pi}(x)$ $=\langle\omega^{(t)},\phi_{\pi}(x)\rangle$
If the Hypothesis 1 holds, then, starting from time $t_{\pi}$ the training dynamics of $\omega^{(t)}$ closely resembles the training dynamics of the kernel machine given by $f^{(t)}_{\pi}$. In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training.
Note 5.1 Hypothesis 1 implies that gradient descent actually learns the implicit kernel machine $f_{\pi}$ and the model $M$ is just a noisy wrapper that allows us to feasibly approximate the extremely-high-dimensional feature function $\phi_{\pi}$.
Since $\pi$ is assumed to be piecewise constant, then also it’s pullback $v_{\pi}$ is piecewise constant and, by Equation 15, we have
$\nabla f_{\pi}(x)=v_{\pi}^{\top}(x)|_{\mathbb{R}^{d_{0}^{*}}}$ (22)
This means that the pullback $v_{\pi}^{\top}$ can actually be interpreted as the gradient of the kernel machine $f_{\pi}$. We can feasibly approximate the direction of that gradient by $v_{\Gamma}$, which is a better approximation than that of $v_{G}$ and therefore $v_{\Gamma}^{\top}$ is actually more faithful to the network’s decision than the standard gradient (if the Hypothesis 1 holds).
Also notice that in order to compute the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$ it suffices to compute the following product kernel:
$K_{\Lambda}(x,x^{\prime}):=\langle x,x^{\prime}\rangle\prod_{\ell=1}^{L}\langle\lambda_{\ell}(x),\lambda_{\ell}(x^{\prime})\rangle=\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle,\quad\text{for }x,x^{\prime}\in\mathbb{R}^{d_{0}}$ (23)
In particular, we may feasibly study the properties of $\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$ with the help of $K_{\Gamma}$ kernel. Notice that $K_{\Gamma}$ encodes the idea behind Deep Features and thus offers a possible explanation for why the latter is an effective perceptual metric (Zhang et al., 2018), as it basically expresses the path-space dot product under the feature map $\phi_{\pi}$ (up to the scaling), tapping into the internal representation of $f_{\pi}$.
5.3 The effectiveness of Batch Normalization
Batch Normalization (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization. While the standard explanation attributes these benefits to reduced internal covariate shift and smoother optimization landscapes (Santurkar et al., 2019), our framework offers an alternative and complementary perspective.
Note that BN basically removes the bias from inputs and largely prevents the emergence of data-dominant output neurons, enforcing their uniform activation across the batch. This facilitates both the neuron specialization and more uniform usage of the layer width. Highly activated neurons are therefore considerably better aligned (directionally) with the particular input then the less activated ones and have larger impact on the weight update during training, which reinforces their role. All these phenomena favor quicker stabilization of path activations, i.e. faster emergence of the path tensor feature and the corresponding linear structure of network representations, improving the training speed and representational quality.
Interestingly, this resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promote weight-input alignment as a mechanism for interpretability. The alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer in standard architectures, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradient, see Section 4.1.
5.4 ReLU network as content-addressable storage
Under the path stability hypothesis, ReLU networks can be interpreted as an associative memory system with excitation pullbacks as its content. Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as binary filtering of atoms. This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s memory space $\mathbb{R}^{\mathcal{P}_{1}}$. During training, for a positive example $x$, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (due to Equation 7), i.e. it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the location $\tilde{G}^{(t)}(x)$. Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$ and therefore $v_{\pi}(x)$ (and $v_{\Gamma}(x)$) contains the most coherent information. This is validated by the perceptual alignment of excitation pullbacks.
During inference, the network first computes the location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored memories $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory. According to the path stability hypothesis, we can extract the actual memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$). It may however change the network’s prediction if the pullbacks weren’t used during the training.